# Logistic Regression makes small LLMs strong and explainable “tens-of-shot” classifiers

### v1

@@ 1.1 Literature review
Instead of learning a linear model on the embeddings, Abbas et al. [2024] used linear probing to calibrate the token probability, an approach we we use as a baseline in our paper as well (PLR-L below).
@@ 2 Performance comparison: Small training set, large test set
We begin in the usual setting in the literature, examining the learning curves as we increase the training set size, with a large pool of samples to draw the training and test set from. Soecifically, we sample 20% of the observations of a data set as the test set and sample training samples of increasing size from the remaining 80% of observations. To obtain stable performance estimates we repeat this procedure 50 times. Confidence intervals show ±$\pm 1.96$ standard errors around the mean performance estimate and are estimated using bootstrapping.
@@
Table 1 shows the performance of PLR-E at different sample sizes. The last two columns show at what sample size (divided by the number of classes) PLR-E outperforms GPT-4. We observe that PLR-E eventually beats GPT-4 in all but four datasets. The last column replicates this analysis but here the embeddings on which we train PLR-E are only based on the sentences, omitting the instructions (see Section 2.2).
@@ 3.1 Validating zero-shot performance
In practice, the user might have some further operational or regulatory reasons for imposing a tight bound on the maximum uncertainty around the observed model performance, which requires a sizeable set of labelled instances. For example, with 25, 50, 100, and 250 labelled data points, the width of the 95% confidence interval around an observed accuracy of 0.75 is 0.36, 0.25, 0.17, and 0.11, respectively.
@@ 5 Explainability
We take an example phrase, $e$, and tokenise it into words and symbols (i.e. not the tokenisation used in an LLM, but in the more conventional meaning). We then calculate the feature importance of the kth token by deleting it and finding the Euclidean distance of the embedding $d_{m}(k,e,c)$ from the classification surface for a given class, $c$ and model $m(t)$, a model with training data size $t$. With $d_{m}(,e,c)$ being the distance for the complete phrase, the feature importance of token $k$ is $f_{m}(k,e,c)=d_{m}(,e,c)-d_{m}(k,e,c)$.
@@ 5.1 Stability and convergence
We take the average of feature importances over a number of independent models with large ($200$ instances) training set size as our baseline, $\tilde{F}^{\infty}(e,c)$. The error for a single model with a smaller training sample size $t$ can be decomposed as $\tilde{F}_{m(t)}(e,c)=[\tilde{F}_{m(t)}(e,c)-\tilde{F}^{t}(e,c)]+[\tilde{F}^{t}(e,c)-\tilde{F}^{\infty}(e,c)]$, where addition and subraction are element-wise in $k$. The first term is the shift in the feature importances for a particular model $m(t)$ trained on a training set of size $t$ versus the average feature importances for all models with a training set of size $t$, and the second term is the shift in that average relative to the average feature importance in the large training set limit.
We report the mean absolute values of both the first and second term in Figure 14. After averaging, the first term describes the spread amongst models for a given training set size and the second to their shift relative to the average over models trained on many training instances ($\tilde{F}^{\infty}(e,c)$). Even at the smallest training sample size the deviations are acceptable for important features which empirically have sizes in $\tilde{F}$ greater than $1$. This can be seen in four random examples in Figure 15 (the results from 16 other examples are shown in Figures C.7C.9 in the appendix). The models are trained on $30$ training instances, which as noted, gives models with a test accuracy on average equal to GPT-4’s. The spreads in the feature importances from models trained on different random $30$-instance training sets are shown by the black bars and are small in fractional terms, particularly for the important features. The shifts relative to the large training set limit, which are not represented in the figures, are around $3$ times smaller than the spreads.
@@ 6 Conclusion
We have have that a penalised regression on embeddings allows local, generative language models to achieve comparable performance to the flagship GPT-4 model in text classification problems. In fact, no more labelled instances are required than are needed for statistically validating the performance of GPT-4. A large number of experiments demonstrated the robustness of our results. An analysis of the embedding space reveals that a handful of the $4096$ embedding vectors often suffice to train an accurate linear model on the embeddings. In addition to general advantages of locally hosted models such as privacy, availability, low cost, we show that our approach enables stable and sensible explanations.
@@ Appendix A Datasets
This dataset contains 2,480 statements of the Federal Open Market Committee based on speeches, meeting minutes and press conferences. The statements have been labelled as having a hawkish, negative or a neutral monetary policy stance. The annotators were given a detailed guide that defined hawkishness and dovishness for eight different categories (economic status, dollar value change, energy/house prices, foreign nations. Fed expectations/actions/assets, money supply, keywords/phrases, labour), which illustrates the complexity of the classification problem. This dataset has been assembled by Shah et al. [2023] and is made available on Github Github. We sample of random subset of 4,419 of the sentences of this dataset.

### v2

@@ 1.1 Literature review
Instead of learning a linear model on the embeddings, Abbas et al. [2024] used linear probing to calibrate the token probability, an approach we use as a baseline in our paper as well (PLR-L below).
@@ 2 Performance comparison: Small training set, large test set
We begin in the usual setting in the literature, examining the learning curves as we increase the training set size, with a large pool of samples to draw the training and test set from. Specifically, we sample 20% of the observations of a data set as the test set and sample training samples of increasing size from the remaining 80% of observations. To obtain stable performance estimates we repeat this procedure 50 times. Confidence intervals show ±$\pm 1.96$ standard errors around the mean performance estimate and are estimated using bootstrapping.
@@
Table 1 shows the performance of PLR-E at different sample sizes. The last two columns show at what sample size (divided by the number of classes) PLR-E outperforms GPT-4. We observe that PLR-E eventually beats GPT-4 in all but four datasets. [^4] Note that the Central Banking data set constitutes a substantially more complex classification problem than the other data sets in our pool. See data set descriptions in Appendix A The last column replicates this analysis but here the embeddings on which we train PLR-E are only based on the sentences, omitting the instructions (see Section 2.2).
@@ 3.1 Validating zero-shot performance
In practice, the user might have some further operational or regulatory reasons for imposing a tight bound on the maximum uncertainty around the observed model performance, which requires a sizeable set of labelled instances. For example, with 25, 50, 100, and 250 labelled data points, the width of the 95% confidence interval around an observed accuracy of 0.75 is 0.36, 0.25, 0.18, and 0.11, respectively.
@@ 5 Explainability
We take an example phrase, $e$, and tokenise it into words and symbols (i.e. not the tokenisation used in an LLM, but in the more conventional meaning). We define feature importance using the decision function ($\boldsymbol{a}.\boldsymbol{e}$ above). We calculate the feature importance of the kth token by deleting it and finding the decision function $d_{m}(k,e,c)$ for a given class, $c$ and model $m(t)$, a model with training data size $t$. Then, with $d_{m}(,e,c)$ being the decision function for the complete phrase, the feature importance of token $k$ is $f_{m}(k,e,c)=d_{m}(,e,c)-d_{m}(k,e,c)$.
@@ 5.1 Stability and convergence
We take the average of feature importances over $20$ independent models with large ($200$ instances) training set size as our baseline, $\tilde{F}^{\infty}(e,c)$. The error for a single model with a smaller training sample size $t$ can be decomposed as $\tilde{F}_{m(t)}(e,c)=[\tilde{F}_{m(t)}(e,c)-\tilde{F}^{t}(e,c)]+[\tilde{F}^{t}(e,c)-\tilde{F}^{\infty}(e,c)]$, where addition and subraction are element-wise in $k$. The first term is the shift in the feature importances for a particular model $m(t)$ trained on a training set of size $t$ versus the average feature importances for all models with a training set of size $t$, and the second term is the shift in that average relative to the average feature importance in the large training set limit.
We report the mean absolute values of both the first and second term in Figure 14. The first term then describes the spread amongst models for a given training set size and the second their shift relative to the average over models trained on many training instances ($\tilde{F}^{\infty}(e,c)$). Even at the smallest training sample size the deviations are acceptable for important features which empirically have sizes in $\tilde{F}$ greater than $1$. This can be seen in four random examples in Figure 15 (the results from 16 other examples are shown in Figures C.7C.9 in the appendix). The models are trained on $30$ training instances, which as noted, gives models with a test accuracy on average equal to GPT-4’s. The spreads in the feature importances from models trained on different random $30$-instance training sets are shown by the black bars and are small in fractional terms, particularly for the important features. The shifts relative to the large training set limit, which are not represented in the figures, are around $3$ times smaller than the spreads.
@@ 6 Conclusion
We have shown that a penalised regression on embeddings allows local, generative language models to achieve comparable performance to the flagship GPT-4 model in text classification problems. In fact, no more labelled instances are required than are needed for statistically validating the performance of GPT-4. A large number of experiments demonstrated the robustness of our results. An analysis of the embedding space reveals that a handful of the $4096$ embedding vectors often suffice to train an accurate linear model on the embeddings. In addition to general advantages of locally hosted models such as privacy, availability, low cost, we show that our approach enables stable and sensible explanations.
@@ Appendix A Datasets
This dataset contains 2,480 statements of the Federal Open Market Committee based on speeches, meeting minutes and press conferences. The statements have been labelled as having a hawkish, negative or a neutral monetary policy stance. The annotators were given a detailed guide that defined hawkishness and dovishness for eight different categories (economic status, dollar value change, energy/house prices, foreign nations. Fed expectations/actions/assets, money supply, keywords/phrases, labour), which illustrates the complexity of the classification problem. This dataset has been assembled by Shah et al. [2023] and is made available on Github Github.